OpenStudy (anonymous):

The difference of the squares of two consecutive odd integers is 48. Find the integers.

OpenStudy (anonymous):

x^2-y^2=48 x-y=2 solve this system.

OpenStudy (anonymous):

how do you solve that? Sorry, I don't find anything in common in order to solve it

OpenStudy (anonymous):

solve for y in the second equation: y=x-2 substitute that into the first equation: x^2 - (x-2)^2 = 48

OpenStudy (anonymous):

now solve that equation involving only x...:)

OpenStudy (anonymous):

Does the negative distribute? @dpaInc

OpenStudy (anonymous):

no.. do the exponentiation first .... expand (x-2)^2 to (x^2-4x+4)

OpenStudy (anonymous):

And so then it becomes -x^2+4x-4? @dpaInc

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

notice the x^2 cancel

OpenStudy (anonymous):

Yup! :) Thank you

OpenStudy (anonymous):

yw..:)